# Generalized Random Forests using Fixed-Point Trees

### v3

@@ B.3 Proofs for Section 4
$=O_{P}\left(r,~{}\frac{1}{\sqrt{n_{C_{1}}}},~{}\frac{1}{\sqrt{n_{C_{2}}}}\right),$
@@
From the proof of Lemma B.4, we know that
$\tilde{\theta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}_{C_{1}}-\tilde{\theta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}_{C_{2}}=\hat{\theta}_{C_{1}}-\hat{\theta}_{C_{2}}+o_{P}\left(r,~{}\frac{1}{\sqrt{n_{C_{1}}}},~{}\frac{1}{\sqrt{n_{C_{2}}}}\right)$
@@
$\tilde{\Delta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2})$ $=\frac{n_{C_{1}}n_{C_{2}}}{n_{P}^{2}}\left\lVert\tilde{\theta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}_{C_{1}}-\tilde{\theta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}_{C_{2}}\right\rVert^{2},$
$=\frac{n_{C_{1}}n_{C_{2}}}{n_{P}^{2}}\left\lVert\hat{\theta}_{C_{1}}-\hat{\theta}_{C_{2}}+o_{P}\left(r,~{}\frac{1}{\sqrt{n_{C_{1}}}},~{}\frac{1}{\sqrt{n_{C_{2}}}}\right)\right\rVert^{2},$
$=\frac{n_{C_{1}}n_{C_{2}}}{n_{P}^{2}}\left\lVert\hat{\theta}_{C_{1}}-\hat{\theta}_{C_{2}}\right\rVert^{2}$
$\hphantom{..........}+\frac{n_{C_{1}}n_{C_{2}}}{n_{P}^{2}}o_{P}\left(r^{2},~{}\frac{1}{n_{C_{1}}},~{}\frac{1}{n_{C_{2}}}\right)$
$\hphantom{..........}+\frac{n_{C_{1}}n_{C_{2}}}{n_{P}^{2}}(\hat{\theta}_{C_{1}}-\hat{\theta}_{C_{2}})o_{P}\left(r,~{}\frac{1}{\sqrt{n_{C_{1}}}},~{}\frac{1}{\sqrt{n_{C_{2}}}}\right),$
$=\frac{n_{C_{1}}n_{C_{2}}}{n_{P}^{2}}\left\lVert\hat{\theta}_{C_{1}}-\hat{\theta}_{C_{2}}\right\rVert^{2}+o_{P}\left(r^{2},~{}\frac{1}{n_{C_{1}}},~{}\frac{1}{n_{C_{2}}}\right)$
$\hphantom{..........}+(\hat{\theta}_{C_{1}}-\hat{\theta}_{C_{2}})o_{P}\left(r,~{}\frac{1}{\sqrt{n_{C_{1}}}},~{}\frac{1}{\sqrt{n_{C_{2}}}}\right)$ (35)
where the final equality holds because $n_{C_{1}}n_{C_{2}}/n_{P}^{2}<1$. Meanwhile, also from Lemma B.4, we know that
$\hat{\theta}_{C_{1}}-\hat{\theta}_{C_{2}}=O_{P}\left(r,~{}\frac{1}{\sqrt{n_{C_{1}}}},~{}\frac{1}{\sqrt{n_{C_{2}}}}\right),$
and hence the cross term (35) can be bound according to the following rate:
$(\hat{\theta}_{C_{1}}-\hat{\theta}_{C_{2}})o_{P}\left(r,~{}\frac{1}{\sqrt{n_{C_{1}}}},~{}\frac{1}{\sqrt{n_{C_{2}}}}\right)$ $=O_{P}\left(r,~{}\frac{1}{\sqrt{n_{C_{1}}}},~{}\frac{1}{\sqrt{n_{C_{2}}}}\right)o_{P}\left(r,~{}\frac{1}{\sqrt{n_{C_{1}}}},~{}\frac{1}{\sqrt{n_{C_{2}}}}\right),$
$=o_{P}\left(r,~{}\frac{1}{\sqrt{n_{C_{1}}}},~{}\frac{1}{\sqrt{n_{C_{2}}}}\right).$
Therefore,
@@
$\tilde{\Delta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2})=\frac{n_{C_{1}}n_{C_{2}}}{n_{P}^{2}}\left\lVert\hat{\theta}_{C_{1}}-\hat{\theta}_{C_{2}}\right\rVert^{2}+o_{P}\left(r^{2},~{}\frac{1}{n_{C_{1}}},~{}\frac{1}{n_{C_{2}}}\right),$
as desired.

### v4

@@ B.2 Proofs for Section 4
$\overline{\psi}_{C_{1}}-\overline{\psi}_{C_{2}}$ $=V(x_{P})\begin{bmatrix}\hat{\theta}_{P}-\theta_{C_{1}}^{*}\\ \hat{\nu}_{P}-\nu_{C_{1}}^{*}\end{bmatrix}-V(x_{P})\begin{bmatrix}\hat{\theta}_{P}-\theta_{C_{2}}^{*}\\ \hat{\nu}_{P}-\nu_{C_{2}}^{*}\end{bmatrix}+O_{P}\left(r^{2},~{}\frac{1}{\sqrt{n_{C_{1}}}},~{}\frac{1}{\sqrt{n_{C_{2}}}}\right),$
$=-V(x_{P})\begin{bmatrix}\theta_{C_{1}}^{*}-\theta_{C_{2}}^{*}\\ \nu_{C_{1}}^{*}-\nu_{C_{2}}^{*}\end{bmatrix}+O_{P}\left(r^{2},~{}\frac{1}{\sqrt{n_{C_{1}}}},~{}\frac{1}{\sqrt{n_{C_{2}}}}\right).$
We assume $\eta$ is a fixed scalar and $\xi^{\top}$ a fixed matrix, it follows:
$\eta\xi^{\top}(\overline{\psi}_{C_{1}}-\overline{\psi}_{C_{2}})=-\eta\xi^{\top}V(x_{P})\begin{bmatrix}\theta_{C_{1}}^{*}-\theta_{C_{2}}^{*}\\ \nu_{C_{1}}^{*}-\nu_{C_{2}}^{*}\end{bmatrix}+O_{P}\left(r^{2},~{}\frac{1}{\sqrt{n_{C_{1}}}},~{}\frac{1}{\sqrt{n_{C_{2}}}}\right).$ (48)
The fixed matrix $\xi^{\top}$ selects the coordinates of the target effect as $\xi^{\top}(\theta,\nu)^{\top}=\theta$, and hence the product $\xi^{\top}V(x_{P})$ simplifies:
$\xi^{\top}V(x_{P})=\xi^{\top}\begin{bmatrix}V_{\theta\theta}(x_{P})&V_{\theta\nu}(x_{P})\\ V_{\nu\theta}(x_{P})&V_{\nu\nu}(x_{P})\end{bmatrix}=\begin{bmatrix}V_{\theta\theta}(x_{P})&V_{\theta\nu}(x_{P})\end{bmatrix}.$
Under Neyman orthogonality one has $V_{\theta\nu}(x_{P})={\bf 0}$, implying that $\xi^{\top}V(x_{P})=[V_{\theta\theta}(x_{P})~{}~{}{\bf 0}]$, and (48) becomes
$\eta\xi^{\top}(\overline{\psi}_{C_{1}}-\overline{\psi}_{C_{2}})=-\eta V_{\theta\theta}(x_{P})(\theta_{C_{1}}^{*}-\theta_{C_{2}}^{*})+O_{P}\left(r^{2},~{}\frac{1}{\sqrt{n_{C_{1}}}},~{}\frac{1}{\sqrt{n_{C_{2}}}}\right).$ (49)
Asymptotic analysis.
Let $E$ be any term satisfying $E=O_{P}(r^{2},~{}1/\sqrt{n_{C_{1}}},~{}1/\sqrt{n_{C_{2}}})$. In our asymptotic regime with $r^{-2}\ll n_{C_{j}}\implies 1/\sqrt{n_{C_{j}}}=o(r)$, one has
$E=O_{P}\left(r^{2},~{}\frac{1}{\sqrt{n_{C_{1}}}},~{}\frac{1}{\sqrt{n_{C_{2}}}}\right)\implies E=o_{P}\left(r,~{}\frac{1}{\sqrt{n_{C_{1}}}},~{}\frac{1}{\sqrt{n_{C_{2}}}}\right)$
and therefore (49) satisfies
$\eta\xi^{\top}(\overline{\psi}_{C_{1}}-\overline{\psi}_{C_{2}})=-\eta V_{\theta\theta}(x_{P})(\theta_{C_{1}}^{*}-\theta_{C_{2}}^{*})+o_{P}\left(r,~{}\frac{1}{\sqrt{n_{C_{1}}}},~{}\frac{1}{\sqrt{n_{C_{2}}}}\right).$ (50)
Write $A=V_{\theta\theta}(\theta_{C_{1}}^{*}-\theta_{C_{2}}^{*})$ such that $\widetilde{\Delta}_{\eta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2})$ is equivalently written $\widetilde{\Delta}_{\eta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2})=(n_{C_{1}}n_{C_{2}}/n_{P}^{2})\cdot\eta^{2}\left\lVert A+E\right\rVert^{2}$. Consider the difference
$\widetilde{\Delta}_{\eta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2})-\frac{n_{C_{1}}n_{C_{2}}}{n_{P}^{2}}\eta^{2}\left\lVert V_{\theta\theta}(\theta_{C_{1}}^{*}-\theta_{C_{2}}^{*})\right\rVert^{2}$ $=\frac{n_{C_{1}}n_{C_{2}}}{n_{P}^{2}}\eta^{2}\left(\left\lVert A+E\right\rVert^{2}-\left\lVert A\right\rVert^{2}\right),$
$=\frac{n_{C_{1}}n_{C_{2}}}{n_{P}^{2}}\eta^{2}\left(2\langle A,E\rangle+\left\lVert E\right\rVert^{2}\right).$
One repeats the same arguments used in the final asymptotic analysis of Lemma B.3 to show
$2\langle A,E\rangle+\left\lVert E\right\rVert^{2}=o_{P}\left(r^{2},~{}\frac{1}{{n_{C_{1}}}},~{}\frac{1}{{n_{C_{2}}}}\right),$
and thus
$\widetilde{\Delta}_{\eta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2})-\frac{n_{C_{1}}n_{C_{2}}}{n_{P}^{2}}\eta^{2}\left\lVert V_{\theta\theta}(x_{P})(\theta_{C_{1}}^{*}-\theta_{C_{2}}^{*})\right\rVert^{2}=o_{P}\left(r^{2},~{}\frac{1}{{n_{C_{1}}}},~{}\frac{1}{{n_{C_{2}}}}\right),$
@@
First, under Assumptions A.1 the $\theta$-block $V_{\theta\theta}(x_{P})$ of the local Jacobian $V(x_{P})$ is strictly positive definite and thus $\left\lVert\cdot\right\rVert_{V}$ defines a true norm. From the proof of Lemma B.3:
$\hat{\theta}_{C_{1}}-\hat{\theta}_{C_{2}}=\theta_{C_{1}}^{*}-\theta_{C_{2}}^{*}+o_{P}\left(r,~{}\frac{1}{\sqrt{n_{C_{1}}}},~{}\frac{1}{\sqrt{n_{C_{2}}}}\right).$
The matrix $V(x_{P})$ is non-random and fixed given $P$ and $\eta$ is a fixed scalar. It follows:
$\eta V_{\theta\theta}(x_{P})(\hat{\theta}_{C_{1}}-\hat{\theta}_{C_{2}})=\eta V_{\theta\theta}(x_{P})(\theta_{C_{1}}^{*}-\theta_{C_{2}}^{*})+o_{P}\left(r,~{}\frac{1}{\sqrt{n_{C_{1}}}},~{}\frac{1}{\sqrt{n_{C_{2}}}}\right).$
Up to a negative factor, the expression on the right is precisely the same as (50) in the proof of Lemma B.4, and thus one repeats the arguments to arrive at
$\left\lVert\eta V_{\theta\theta}(x_{P})(\hat{\theta}_{C_{1}}-\hat{\theta}_{C_{2}})\right\rVert_{2}^{2}=\left\lVert\eta V_{\theta\theta}(x_{P})(\theta_{C_{1}}^{*}-\theta_{C_{2}}^{*})\right\rVert_{2}^{2}+o_{P}\left(r^{2},~{}\frac{1}{{n_{C_{1}}}},~{}\frac{1}{{n_{C_{2}}}}\right),$
@@
$\Delta_{\eta V}(C_{1},C_{2})=\frac{n_{C_{1}}n_{C_{2}}}{n_{P}^{2}}\eta^{2}\left\lVert V_{\theta\theta}(x_{P})(\theta_{C_{1}}^{*}-\theta_{C_{2}}^{*})\right\rVert_{2}^{2}+o_{P}\left(r^{2},~{}\frac{1}{{n_{C_{1}}}},~{}\frac{1}{{n_{C_{2}}}}\right).$
@@
The right hand side is precisely the same as in the statement of Lemma B.4 established for $\widetilde{\Delta}_{\eta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2})$, and thus
@@
$\widetilde{\Delta}_{\eta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2})-\Delta_{\eta V}(C_{1},C_{2})=o_{P}\left(r^{2},~{}\frac{1}{{n_{C_{1}}}},~{}\frac{1}{{n_{C_{2}}}}\right),$ (51)
as desired. ∎
Proof of Lemma 4.2.
Firstly, Specifications 4 (subsampling) and 5 (honesty) describe conditions imposed on the sampling mechanism and are not affected by the form of the splitting criterion. It remains to verify whether Specification 1 (symmetry), Specification 2 (balanced/$\omega$-regular), and Specification 3 (randomized/random-split) are satisfied by $\mathcal{T}(\Delta_{V})$.
  1. 1.
    Specification 1: Symmetry. A tree is said to be symmetric if its estimates are invariant under permutations of the tree’s training samples. Conditional on a sequence of criterion values computed over splits of $P$, the CART mechanism of selecting the best split by scanning over the collection of candidates does not depend on the parent samples at all. This means that asymmetry could only enter through the criterion values. Therefore, a sufficient condition for symmetry in the tree estimates with respect to permutations of the tree samples is whether the criterion $\Delta_{V}(C_{1},C_{2})$ is symmetric. Conditional on the child solutions $\hat{\theta}_{C_{1}},\hat{\theta}_{C_{2}}$, the map $(\hat{\theta}_{C_{1}},\hat{\theta}_{C_{2}})\mapsto\Delta_{V}(C_{1},C_{2})$ does not depend on the parent samples at all, and therefore asymmetry could only enter through child solutions $\hat{\theta}_{C_{j}}$. However, both criteria use precisely the same child solutions $\hat{\theta}_{C_{j}}$ in (7), and therefore $\Delta_{V}(C_{1},C_{2})$ will be symmetric whenever $\Delta(C_{1},C_{2})$ is symmetric (specifically, whenever $\psi_{\theta,\nu}(O_{i})$ is symmetric with respect to permutations of $i$).
  2. 2.
    Specification 2: Balanced/$\omega$-regular. This condition is enforced by by GRF by adding an additional stopping condition to the gradient-based version of Algorithm 1. Specifically, GRF stops a recursive splitting path if a proposed $\Delta$-optimal split were to send fewer than $\omega n_{P}$ of the parent samples into either child. Simply stated, GRF enforces balanced splits by defining the set of valid candidate splits as those with at least $\omega n_{P}$ parent samples in each child. This is left unchanged by our method.
  3. 3.
    Specification 3: Randomized/random-split. The asymptotic theory of GRF requires that, at each node, each variable is selected for a split with some lower bound probability $\pi>0$. In order to satisfy the minimum split probability GRF uses the feature sampling device of Denil et al. (2014) which, at each step, considers only $\min\{\max\{\textrm{Poisson}(m),1\},p\}$ randomly selected features as candidate variables, where $p=\dim(\mathcal{X})$ and $m$ is a GRF tuning parameter. In other words, GRF defines the set of valid candidate splits such that the set of valid splitting dimensions is itself a random variable. This mechanism is left unchanged under our method.
    No column of $V(x_{P})$ is all-zero $V_{\cdot,k}(x_{P})\neq{\bf 0}$ because $V(x_{P})$ is strictly positive definite symmetric, and therefore $\Delta_{V}(C_{1},C_{2})$ will not systematically ignore signals along parameter dimensions $\theta_{k}$ that can be detected by the $\Delta$-criterion. Finally,
    $\hat{\theta}_{C_{1}}-\hat{\theta}_{C_{1}}\neq{\bf 0}\implies V_{\theta\theta}(x_{P})(\hat{\theta}_{C_{1}}-\hat{\theta}_{C_{2}})\neq{\bf 0},$
    because $V(x)$ is strictly positive definite symmetric by Assumption 2. Therefore $\Delta(C_{1},C_{2})>0\implies\Delta_{V}(C_{1},C_{2})>0$ meaning that the $\Delta_{V}$-criterion is non-degenerate and will always be able to select at least one feature whenever the $\Delta$-criterion can select a feature.
Therefore, all five specifications are met, and one concludes that $\mathcal{T}(\Delta_{V})$ must satisfy the forest Specifications A.2 whenever they are satisfied by $\mathcal{T}(\Delta)$. ∎
B.3 Asymptotic equivalence of the pseudo-outcome approximation for VCM/HTE models
In this section we establish the asymptotic equivalence of the further acceleration of the fixed-point method proposed in Section 5 for VCM/HTE models. The accelerated algorithm is based on FPT pseudo-outcomes that use an approximation $\tilde{\theta}_{P}$ for the actual parent solutions $\hat{\theta}_{P}$ in (31). Specifically, the parent-leaf approximations $\tilde{\theta}_{P}$ are found by a single gradient descent step towards $\hat{\theta}_{P}$ taken from the origin (54). Let $\rho_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$ denote the original FPT pseudo-outcomes for VCM/HTE models:
$\rho_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}\coloneqq-(W_{i}-\overline{W}_{P})\left(Y_{i}-\overline{Y}_{P}-(W_{i}-\overline{W}_{P})^{\top}\hat{\theta}_{P}\right),$ (52)
where the solution $\hat{\theta}_{P}$ for the local model over the parent $P$ are precisely the OLS coefficients from the regression the centered outcomes $Y_{i}-\overline{Y}_{P}\in\mathbb{R}$ on the centered regressors $W_{i}-\overline{W}_{P}\in\mathbb{R}^{K}$. In contrast, let $\phi_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$ denote approximations of $\rho_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$ pseudo-outcomes that are of the form
$\phi_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}\coloneqq-(W_{i}-\overline{W}_{P})\left(Y_{i}-\overline{Y}_{P}-(W_{i}-\overline{W}_{P})^{\top}\tilde{\theta}_{P}\right),$ (53)
where $\tilde{\theta}_{P}$ approximates $\hat{\theta}_{P}$ as:
$\tilde{\theta}_{P}\coloneqq\gamma\cdot\frac{1}{n_{P}}\sum_{\{i:X_{i}\in P\}}(W_{i}-\overline{W}_{P})(Y_{i}-\overline{Y}_{P})=\gamma\cdot\frac{1}{n_{P}}W_{P}^{\top}Y_{P}.$ (54)
Here, $W_{P}\in\mathbb{R}^{n_{P}\times K}$ and $Y_{P}\in\mathbb{R}^{n_{P}}$ denote the centered data matrices, $W_{P}\coloneqq[W_{i}-\overline{W}_{P}]_{i:X_{i}\in P}$ and $Y_{P}\coloneqq[Y_{i}-\overline{Y}_{P}]_{i:X_{i}\in P}$, and the scalar $\gamma>0$ denotes the exact line search step size corresponding to the regression of the centered outcomes on the centered regressors:
$\gamma\coloneqq\frac{\left\lVert W_{P}^{\top}Y_{P}\right\rVert_{2}^{2}}{\left\lVert W_{P}W_{P}^{\top}Y_{P}\right\rVert_{2}^{2}}.$ (55)
Lemma B.5.
Let $\tilde{\theta}_{C_{j}}$ denote the FPT estimator of the form (24) for the child solution $\hat{\theta}_{C_{j}}$ for VCM/HTE models. One can express $\tilde{\theta}_{C_{j}}$ in terms of the corresponding fixed-point pseudo-outcomes:
$\tilde{\theta}_{C_{j}}\coloneqq\hat{\theta}_{P}+\frac{1}{n_{C_{j}}}\sum_{i:X_{i}\in C_{j}}\rho_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}.$
Similarly, denote by $\bar{\theta}_{C_{j}}$ the FPT estimator of $\hat{\theta}_{C_{j}}$ induced by pseudo-outcomes approximations $\phi_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$:
$\bar{\theta}_{C_{j}}\coloneqq\hat{\theta}_{P}+\frac{1}{n_{C_{j}}}\sum_{\{i:X_{i}\in C_{j}\}}\phi_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}.$
Then, under the assumptions of Proposition 4.1, $\bar{\theta}_{C_{j}}$ is consistent for $\tilde{\theta}_{C_{j}}$ as:
$\left\lVert\tilde{\theta}_{C_{j}}-\bar{\theta}_{C_{j}}\right\rVert=o_{P}(1).$
Proof.
A direct calculation reveals that the difference between the original FPT pseudo-outcomes $\rho_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$ in (52) and the approximations $\phi_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$ in (53) satisfy
$\rho_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}-\phi_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$ $=-(W_{i}-\overline{W}_{P})\left(\left[Y_{i}-\overline{Y}_{P}-(W_{i}-\overline{W}_{P})^{\top}\hat{\theta}_{P}\right]-\left[Y_{i}-\overline{Y}_{P}-(W_{i}-\overline{W}_{P})^{\top}\tilde{\theta}_{P}\right]\right),$
$=(W_{i}-\overline{W}_{P})(W_{i}-\overline{W}_{P})^{\top}(\hat{\theta}_{P}-\tilde{\theta}_{P}).$ (56)
Therefore, the difference between the original FPT child estimator $\tilde{\theta}_{C_{j}}$ and the approximation $\bar{\theta}_{C_{j}}$ satisfies
$\tilde{\theta}_{C_{j}}-\bar{\theta}_{C_{j}}$ $=\frac{1}{n_{C_{j}}}\sum_{\{i:X_{i}\in C_{j}\}}(\rho_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}-\phi_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}),$
$=\frac{1}{n_{C_{j}}}\sum_{\{i:X_{i}\in C_{j}\}}(W_{i}-\overline{W}_{P})(W_{i}-\overline{W}_{P})^{\top}(\hat{\theta}_{P}-\tilde{\theta}_{P}),$
$=S_{C_{j}}(\hat{\theta}_{P}-\tilde{\theta}_{P}),$
where we denote $S_{C_{j}}\coloneqq\frac{1}{n_{C_{j}}}\sum_{\{i:X_{i}\in C_{j}\}}(W_{i}-\overline{W}_{P})(W_{i}-\overline{W}_{P})^{\top}$. Therefore,
$\left\lVert\tilde{\theta}_{C_{j}}-\bar{\theta}_{C_{j}}\right\rVert=\left\lVert S_{C_{j}}(\hat{\theta}_{P}-\tilde{\theta}_{P})\right\rVert\leq\left\lVert S_{C_{j}}\right\rVert_{F}\left\lVert\hat{\theta}_{P}-\tilde{\theta}_{P}\right\rVert.$ (57)
Under GRF’s regularity conditions, in a limit where $n_{C_{j}}\to\infty$ and the parent radius $r\coloneqq\sup_{\{i:X_{i}\in P\}}\left\lVert X_{i}-\overline{X}_{P}\right\rVert$ goes to zero $r\to 0$, we have $S_{C_{j}}\stackrel{{\scriptstyle p}}{{\to}}Q$ for some positive semidefinite symmetric matrix $Q$, and hence $\left\lVert S_{C_{j}}\right\rVert_{F}=O_{P}(1)$. Meanwhile, by definition (31) for the OLS coefficients $\hat{\theta}_{P}$ and definition (54) for the one-step approximations $\bar{\theta}_{P}$, we have
$\left\lVert\hat{\theta}_{P}-\tilde{\theta}_{P}\right\rVert$ $=\left\lVert\left(-A_{P}^{-1}\cdot n_{P}^{-1}W_{P}^{\top}Y_{P}\right)-\left(\gamma\cdot n_{P}^{-1}W_{P}^{\top}Y_{P}\right)\right\rVert,$
$=\left\lVert\left(-A_{P}^{-1}-\gamma\mathbb{I}\right)\cdot n_{P}^{-1}W_{P}^{\top}Y_{P}\right\rVert,$
$\leq\left\lVert-A_{P}^{-1}-\gamma\mathbb{I}\right\rVert_{F}\left\lVert n_{P}^{-1}W_{P}^{\top}Y_{P}\right\rVert,$
$=\left\lVert\left[n_{P}^{-1}W_{P}^{\top}W_{P}\right]^{-1}-\gamma\mathbb{I}\right\rVert_{F}\left\lVert n_{P}^{-1}W_{P}^{\top}Y_{P}\right\rVert,$ (58)
where $-A_{P}=n_{P}^{-1}W_{P}^{\top}W_{P}$ follows from the definition of $A_{P}$ as an estimator of the Jacobian $\nabla\psi$, e.g. (13) in the context of VCM/HTE models. Under the Lipschitz continuity Assumptions 13, one has the standard stochastic bound for the cross term $n_{P}^{-1}W_{P}^{\top}Y_{P}$:
$\left\lVert n_{P}^{-1}W_{P}^{\top}Y_{P}\right\rVert=O_{P}\left(r,~{}\frac{1}{\sqrt{n_{P}}}\right),$ (59)
while the difference $[n_{P}^{-1}W_{P}^{\top}W_{P}]^{-1}-\gamma\mathbb{I}$ is stochastically bound as
$\left\lVert\left[n_{P}^{-1}W_{P}^{\top}W_{P}\right]^{-1}-\gamma\mathbb{I}\right\rVert_{F}=O_{P}(1),$
because $n_{P}^{-1}W_{P}^{\top}W_{P}\stackrel{{\scriptstyle p}}{{\to}}\text{Cov}(W_{i}\mid X_{i}\in P)$ is non-singular under Assumption 2. Coupling these stochastic bounds together according to (58) gives
$\left\lVert\hat{\theta}_{P}-\tilde{\theta}_{P}\right\rVert=O_{P}\left(r,~{}\frac{1}{\sqrt{n_{P}}}\right),$
and trivially, because $n_{C_{j}}<n_{P}$,
$\left\lVert\hat{\theta}_{P}-\tilde{\theta}_{P}\right\rVert=O_{P}\left(r,~{}\frac{1}{\sqrt{n_{C_{j}}}}\right).$ (60)
Under Proposition 1 of GRF one assumes $r^{-2}\ll n_{C_{1}},n_{C_{2}}$ and thus, in an asymptotic regime where $n_{C_{j}}\to\infty$ and $r\to 0$, one has $1/\sqrt{n_{C_{j}}}=o(r)$, and hence:
$\left\lVert\hat{\theta}_{P}-\tilde{\theta}_{P}\right\rVert=o_{P}(1).$ (61)
Returning to (57), the consistency of the parent approximation $\tilde{\theta}_{P}$ as (61) implies that the approximation $\bar{\theta}_{C_{j}}$ is itself consistent for the original FPT child estimator $\tilde{\theta}_{C_{j}}$:
$\left\lVert\tilde{\theta}_{C_{j}}-\bar{\theta}_{C_{j}}\right\rVert=o_{P}(1),$ (62)
as desired. ∎